OpenStudy (anonymous):

Let A=|-39 28| |-60 43| has an eigenvalue of lambda = 1. An eigenvector for lambda = 1 has the form | a | |-10| What is the value of a? (If the values in A are too large or unwieldy, you may be able to regenerate a new problem depending on the settings of this assignment.)

OpenStudy (anonymous):

let the eigenvector we are finding for be x. now we know that: Ax=lambda*x Ax=x so what we need to do is solve with this matrix: \[\left[\begin{matrix}-39 & 28 \\ -60 & 43\end{matrix}\right]\left(\begin{matrix}a \\ -10\end{matrix}\right)=\left(\begin{matrix}a \\ -10\end{matrix}\right)\] now to find a, we see that we only need to solve this simple: -39a-280=a